# The Wisdom of a Crowd of Brains: A Universal Brain Encoder

### v1

@@ Appendix A Further Technical Details
The datasets utilized in our study comprise BOLD fMRI responses to various natural images, recorded over multiple scanning sessions. Peak BOLD responses corresponding to each stimulus were estimated. Each dataset underwent unique pre-processing procedures, detailed in their respective publications [23, 39, 44, 43].
@@
For the NSD dataset [43], we used a post-processed dataset with voxel selection from [45]. They chose vision-related areas, resulting in a total of approximately 40,000 voxels, and provided voxels after per voxel normalization.
“fMRI-on-ImageNet" Dataset.
For the fMRI-on-ImageNet dataset [41], a relevant set of around 5,000 voxels was already provided. We implemented Z-scoring normalization ourselves for this dataset.
@@
For the VIM1 dataset [23, 39], we selected the best 7,000 voxels according to the highest Signal-to-Noise Ratio (SNR). SNR is calculated as the ratio of the variance of averaged (repeated) measurements for different stimuli to the average variance of measurements for the same stimuli. This approach ensures the selection of voxels most representative of neural activity in response to diverse visual stimuli. Z-scoring normalization was also implemented for this dataset.
A.2 modified LoRa
The Lora adaptation  [38] is done by adding learned low rank matrices to weights in the original network. In the original paper the value projection and query projection weights ($W_{v}$,$W_{q}$) in the self attention block are modified. We only modify the output projection weights($W_{o}$) of the self attention block.
@@ Appendix B Limitations
Appendix C Additional Figures
Figure S8: Retrieval Results: Our universal encoder trained on multiple subjects significantly improves retrieval accuracy over both subject-specific models as evidenced by the calculated p-value. We conducted 10 retrieval experiments, wherein we randomly sampled 999 different distracted images for each “query” fMRI. Results are compared across three models: (i) "Baseline Encoder" – the encoder of [15] trained separately on each subject, (ii) "Universal Encoder - Single Subject" – our architecture trained separately on each subject, and (iii) "Universal Encoder - Multiple Subjects" – our model trained on combined data from 8 subjects. In (a) we present the mean top-1 accuracy (across the 10 experiments) along with the standard deviation for each model and subject. (b) depicts the mean top-5 accuracy along with the standard deviation for each model and subject. (c) Table of p-values per subject illustrates the improvement significance of our universal encoder trained on multiple subjects compared to the other two models, emphasizing the significance of the improvement.
@@ Appendix C Additional Figures
Figure S9: Transfer Learning Significance Results: The transferred Universal-Encoder significantly outperforms any single-subject model on all datasets. This figure shows the p-value of the median Pearson correlation between the predicted and real fMRI scan for each dataset across 5 runs. For each dataset we compared: (i) Transfer-Learning of the pre-trained Universal-Encoder to the new subject, with varying numbers of subject-specific training data, and (ii) a dedicated subject-specific model, trained from scratch on the subject-specific data only. This extends the findings presented in Fig. 6.
@@
Figure S10: Voxel Embedding Clusters: 5-top images which activated voxels within each cluster. Most clusters are consistent across subjects indicating that voxel embeddings capture functional roles rather than individual identities.

### v2

@@ Appendix A Further Technical Details
The datasets utilized in our study comprise BOLD fMRI responses to various natural images, recorded over multiple scanning sessions. Peak BOLD responses corresponding to each stimulus were estimated, resulting in a scalar value for each voxel for each image. Each dataset underwent unique pre-processing procedures, detailed in their respective publications (Kay et al., 2008; Naselaris et al., 2009; Horikawa & Kamitani, 2017a; Allen et al., 2022).
@@
For the NSD dataset (Allen et al., 2022), we used a post-processed dataset with voxel selection from (Gifford et al., 2023). They chose vision-related areas, resulting in a total of approximately 40,000 voxels, and provided voxels after per voxel normalization.
“fMRI-on-ImageNet” Dataset.
For the fMRI-on-ImageNet dataset (Horikawa & Kamitani, 2017b), a relevant set of around 5,000 voxels was already provided. We implemented Z-scoring normalization ourselves for this dataset.
@@
For the VIM1 dataset (Kay et al., 2008; Naselaris et al., 2009), we selected the best 7,000 voxels according to the highest Signal-to-Noise Ratio (SNR). SNR is calculated as the ratio of the variance of averaged (repeated) measurements for different stimuli to the average variance of measurements for the same stimuli. This approach ensures the selection of voxels most representative of neural activity in response to diverse visual stimuli. Z-scoring normalization was also implemented for this dataset.
A.2 Training and Lora Modification
The Lora adaptation  (Hu et al., 2021) is done by adding learned low rank matrices to weights in the original network. In the original paper the value projection and query projection weights ($W_{v}$,$W_{q}$) in the self attention block are modified. We only modify the output projection weights($W_{o}$) of the self attention block.
A.3 Training
The model is trained end-to-end (Adam optimizer with learning rate of 1e-3). The model’s objective is to correctly predict the voxel fMRI activation on each input image. Each training batch contains 32 randomly selected images (we randomly sample 5000 voxel indices per image, with their corresponding fMRI activations). Each subject (brain) has its own unique voxel indices. Our loss function is: $\mathcal{L}\left(\hat{r},r\right)=\alpha\cdot\text{MSE}\left(\hat{r},r\right)-\left(1-\alpha\right)\cos\left(\angle\left(\hat{r},r\right)\right),$ where $\hat{r}$ and $r$ are the predicted and measured fMRI activations, respectively, and $\alpha$=0.1. Training the Universal-Encoder jointly on 8 NSD subjects (see “Datasets” below), takes $\sim$1 day on a single Quadro RTX 8000 GPU. Inference time (Image-to-fMRI encoding) takes $\sim$50 ms per-image.
@@ Appendix B Limitations
Appendix C Statistical Tests
We conducted statistical tests to evaluate the significance of the Universal Encoder’s performance across the various metrics assessed.
To evaluate the statistical significance of the prediction correlation gap between our multi-subject Universal Encoder’s and the other two models as reported in 4.1 and shown in Fig. 4a , we performed a repeated measures ANOVA test. The analysis, based on the predicted voxel correlations of the 8 NSD subjects, revealed a significant performance gap between models: F-value of 72.74 and p-value of 0 (Num DF=2, Den DF=14).
Additionally, we tested the prediction correlations of individual voxels (across all subjects) against a null hypothesis. The null hypothesis was based on the distribution of correlations between pairs of independent Gaussian random vectors of size N = 982, matching the number of image-fMRI pairs in the test set. The one-sided statistical significance was estimated by comparing the predicted voxel correlations for each subject on the test set with those from the null distribution. We applied a statistical threshold of P $<$ 0.05, correcting for multiple comparisons using the FDR procedure. Our results indicate that the majority of voxel predictions by the Universal Encoder are statistically significant, with an average of 98% across all subjects. Detailed results in Table S1.
Subj1 Subj2 Subj3 Subj4 Subj5 Subj6 Subj7 Subj8
Total Voxels 39,548 39,548 39,548 39,548 39,548 39,198 39,198 39,511
Significantly Predicted Voxel 39,299 39,024 38,612 39,059 39,252 38,665 38,754 36,850
Significant Voxel Percentage 99.37% 98.68% 97.63% 98.77% 99.25% 98.64% 98.87% 93.26%
Table S1: Significance Test Comparison with Null Distribution. Number and percentage of statistically significant voxels for each subject.
Furthermore, we present the significance of the retrieval results reported in 4.1 and shown in Fig. 4b. We performed a t-test by repeating the retrieval experiment 10 times with random image distractors and comparing the average retrieval performance per repeat of our multi-subject Universal Encoder against a competing model. The resulting p-values are presented in Table S2, where the highest p-value observed is approximately $\sim 9e-13$. This illustrates the significance of the Universal Encoder’s improvements when trained on multiple subjects, compared to the other models. The results were corrected for multiple comparisons using the FDR procedure.
Lastly, we present the the significance of the transfer learning results reported in 4.3 and shown in Figure 6. We performed a t-test on the median Pearson correlation between predicted and real fMRI scans, comparing transfer learning with our universal encoder against baseline model, over 5 runs with different random model initializations, this was done for each of the dataset. Table S3 contains the estimated p-values. The results were corrected for multiple comparisons using the FDR procedure.
Subject Universal encoder - multiple subjects Universal encoder - multiple subjects
Vs Universal encoder - single subject Vs Baseline encoder
1 3.795e-14 1.284e-23
2 9.275e-13 8.783e-16
3 9.848e-19 2.831e-25
4 5.274e-19 3.413e-22
5 1.407e-15 6.501e-22
6 3.413e-22 1.337e-23
7 7.104e-16 3.633e-21
8 8.491e-21 1.824e-21
Table S2: Retrieval Results: Our universal encoder trained on multiple subjects significantly improves retrieval accuracy over both subject-specific models as evidenced by the calculated t-test results. t-test was conducted by repeating the experiment 10 times with random image distractors and comparing the average retrieval performance for each repeat. Table of p-values per subject illustrates the improvement significance of our universal encoder trained on multiple subjects compared to the other two models, emphasizing the significance of the improvement. The results were corrected for multiple comparisons using the FDR procedure.
Samples NSD fMRI-on-ImageNet VIM1
100 4.85e-08 2.27e-05 1.66e-01
200 2.69e-06 4.12e-07 1.15e-02
400 1.68e-08 7.42e-05 2.56e-04
600 1.49e-06 9.32e-07
800 1.87e-05 9.22e-05 4.89e-06
1000 2.38e-05 1.39e-04
1200 5.18e-07
1250 3.97e-05
1500 1.07e-04
1600 3.64e-07
2400 2.69e-06
3200 2.41e-08
6400 2.01e-05
@@ Appendix C Statistical Tests
Table S3: Transfer Learning Significance Results: The transferred Universal-Encoder significantly outperforms any single-subject model on all datasets. This figure shows t-test’s p-values of the median Pearson correlation between the predicted and real fMRI scan for each dataset across 5 runs with different random model initialization. For each dataset we compared: (i) Transfer-Learning of the pre-trained Universal-Encoder to the new subject, with varying numbers of subject-specific training data, and (ii) a dedicated subject-specific model, trained from scratch on the subject-specific data only. This extends the findings presented in Fig. 6. The results were corrected for multiple comparisons using the FDR procedure.
Appendix D Additional Figures
Figure S8: Comparison against additional Encoder models. Pearson correlation between predicted and ground-truth fMRI of Subject1 from NSD dataset. The figure presents comparisons across three prominent encoders from different papers (Gaziv et al., 2022; Wang et al., 2023; Takagi & Nishimoto, 2023), as well as the universal encoder trained on both single-subject and multi-subject data. (a) Shows the median values along with the 25th and 75th percentiles across the five models. (b) Compares correlation scores across different coarse brain regions (ROIs) for the same five models. .
Figure S9: Ablation of the Universal Encoder components We compare our full Universal Encoder model (green) to three ablated versions, all trained on data from 8 NSD subjects and evaluated on each subject’s test set. (i) A shared latent space approach, where the cross-attention block is replaced with a projection to a shared latent space for all subjects (blue). (ii) Our model without the spatial attention component (red). (iii) Our model without the functional attention component (yellow). (a) Pearson correlation (per voxel) between predicted & ground-truth fMRI (Median value). (b) Retrieval accuracy: percentage of times the GT image being the Top-1 retrieved image (out of 1000) using the ”Query” fMRI. As can be seen the shared latent space approach yields significantly poorer results. Furthermore, as evidenced by both the retrieval and correlation results, each component (functional and spatial) contributes substantially to performance improvement over the shared space approach, with the combination of both providing the best results.
Figure S10: Where on the brain is the error different for the different models. (a) Higher granularity of brain regions (ROIs), compared across the best 3 models for subject 1. (b) A visual display of the voxel correlation maps on top of the brain shown for 2 models for subject 1 – the Universal Encoder multiple-subjects versus the Baseline Encoder.
Figure S11: Retrieval Results: Our universal encoder trained on multiple subjects significantly improves retrieval accuracy over both subject-specific models as evidenced by the calculated p-value. We conducted 10 retrieval experiments, wherein we randomly sampled 999 different distracted images for each “query” fMRI. Results are compared across three models: (i) ”Baseline Encoder” – the encoder of (Gaziv et al., 2022) trained separately on each subject, (ii) ”Universal Encoder - Single Subject” – our architecture trained separately on each subject, and (iii) ”Universal Encoder - Multiple Subjects” – our model trained on combined data from 8 subjects. In (a) we present the mean top-1 accuracy (across the 10 experiments) along with the standard deviation for each model and subject. (b) depicts the mean top-5 accuracy along with the standard deviation for each model and subject.
Figure S12: MSE Evaluation - The Wisdom of a Crowd of Brains. By aggregating data from multiple subjects, our Universal-Encoder improves the encoding of any individual subject. We present results calculating the MSE between predicted and ground-truth fMRI (median value with 25th and 75th percentiles). We compared 3 models: (i) The ”Baseline” single-subject encoder of (Gaziv et al., 2022), (ii) ”Universal Encoder - single subject” – our architecture trained on each subject separately, (iii) ”Universal Encoder - multiple subjects” – our model trained on data from 8 subjects.
Voxel Embedding Dimension 64 128 256 512 1024
Median Pearson Correlation 0.5170 0.5221 0.5221 0.5227 0.5243
Table S4: Ablation on Different Voxel Embedding Dimensions. Ablation study on the influence of voxel embedding dimensions E on the Universal Encoder’s performance, evaluated on Subject 1 from the NSD dataset. The results show minimal sensitivity to E, indicating that encoding performance remains consistent across different embedding sizes.
Figure S13: Qualitative Evaluation of Image Reconstruction from fMRI. Visual comparison of image reconstruction from fMRI for Subject 1 from the NSD dataset. The figure showcases the ground truth (GT) image seen by the subject, alongside reconstructions generated using the fMRI-to-Image decoding model (MindEye (Scotti et al., 2024a)) trained on Subject 1’s fMRI data. Both the decoder and our Universal Encoder were trained on the same dataset and evaluated on the same test set. Reconstructions were evaluated on the test set using two inputs: (1) ”Reconstruction from GT fMRI” where the decoder processes the real fMRI recordings of the test images, and (2) ”Reconstruction from our encoded fMRI”, where the test images are first encoded into predicted fMRI using our Universal Encoder before being decoded. Both successful reconstructions and failure cases are presented. As observed, the reconstructions from the encoded fMRI are on par with those from the GT fMRI, despite the decoder being trained solely on GT fMRI. This highlights the effectiveness of our Universal Encoder, demonstrating that the encoded fMRI retains information exists in the GT fMRI.
MAE $\downarrow$ MSE $\downarrow$ Lpips $\downarrow$ SSIM $\uparrow$
Reconstruction from GT fMRI 4.063 26.965 0.712 0.139
Reconstruction from our encoded fMRI 4.089 27.002 0.710 0.123
Table S5: Quantitative Evaluation of Image Reconstruction from fMRI. Quantitative evaluation of image reconstruction quality comparing the ground truth (GT) images to: (i) reconstructions from GT fMRI and (ii) reconstructions from encoded fMRI. The results presented represent the metrics averaged over approximately 1,000 images from Subject 1’s test data, which serves as the test set for both the encoder and decoder. Metrics include MAE (Mean Absolute Error), MSE (Mean Squared Error), LPIPS (Learned Perceptual Image Patch Similarity, where lower values indicate closer perceptual similarity), and SSIM (Structural Similarity Index Measure, where higher values indicate greater structural similarity). Results demonstrate that reconstructions from encoded fMRI are comparable to those from GT fMRI, further validating the effectiveness of our Universal Encoder in preserving critical image-relevant fMRI features.
Figure S14: Voxel Embedding Clusters: These are 20 clusters obtained by applying k-means to the voxel embeddings from the left and right hemispheres of subjects 1 and 2, with each cluster represented in a different color. As discussed in the paper, we further investigate the functional role of each cluster.
@@ Appendix D Additional Figures
Figure S15: Voxel Embedding Clusters: Top images which activated voxels within each cluster. Most clusters are consistent across subjects (marked with red frame) indicating that voxel embeddings capture functional roles rather than individual identities.
@@
Figure S17: 2D t-SNE Visualization of Brain Streams. This t-SNE visualization represents voxel embeddings (50,000 voxels were randomly sampled) from all 8 subjects (from NSD), with colors indicating different brain streams (e.g., early, lateral, and parietal streams) and shapes representing different subjects. The separation of brain streams in the voxel-embedding space demonstrates that our embedding effectively captures stream-specific functional properties, while embeddings from different subjects intermingle, reflecting shared functionality across subjects.
Figure S18: 2D t-SNE Visualization of Category-Selective Regions. This t-SNE visualization highlights embeddings from four predefined category-selective brain regions: faces (e.g., FFA, OFA), bodies (e.g., EBA, FBA), places (e.g., PPA, OPA, RSC), and words (e.g., VWFA, OWFA, mfs-words). The colors signify the pre-defined cortical regions, whereas the proximity is imposed by the similarity of our voxel-embeddings. For visualization we randomly chose 50,000 voxels. As can be seen, our voxel embedding implicitly learned functionality which is similar to the functionality of known category selective brain regions.